OpenStudy (anonymous):

Solve for x:5/2 x +15/2 = 15

OpenStudy (anonymous):

combine like terms using additions or subtraction.

OpenStudy (anonymous):

2.5x+7.5=15 -7.5 -7.5 2.5x=7.5 2.5x/2.5=7.5x/2.5 x=3 Answer: x=3

OpenStudy (anonymous):

X=3

OpenStudy (anonymous):

since 15/12 is positive then you would subtract 15/2 from 15 and get... then divide that answer with x